# LayerPlexRank: Exploring Node Centrality and Layer Influence through Algebraic Connectivity in Multiplex Networks

### v1

@@ Abstract.
Centrality, Influence, Algebraic Connectivity, Multiplex Networks
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Mathematics of computing Graph algorithmsccs: Mathematics of computing Network flowsccs: Information systems Content rankingccs: Information systems Learning to rankccs: Information systems Similarity measuresccs: Theory of computation Network flows
@@ 3. Experiments and Evaluation
Refer to captionRefer to caption
Refer to captionRefer to caption
@@ 3.3. Robustness by Designed Cross-validation for Multiplex Networks
In experiments, we demonstrate robustness using a specifically designed Leave-one-out Cross Validation (LOOCV) method (Sammut and Webb, 2010). LOOCV can be considered a specific application of cross-validation, which splits the original $n$-node dataset into an $n-1$ node training dataset and a $1$-node validation dataset in each iteration. The total number of validation times is $n$, indicating that all nodes have been left out once and used as a validation dataset. However, LOOCV focuses on multiple evaluations of the remaining training set using the split validation set, while our algorithm does not predict unknown data from existing data.
Therefore, we incorporate Jackknife resampling (Berger and Rao, 2006) to enhance this validation, which intentionally omits a node and its connected edges in each iteration, combining well with LOOCV to robustly aggregate resampling results. This method underscores the significant resistance of our algorithm against disruptions.
@@ 3.4. Parameter Sensitivity Analysis
Refer to captionRefer to caption
Refer to captionRefer to caption
@@ 4. Conclusions
LayerPlexRank significantly advances the field of multiplex network analysis by integrating node centrality and layer influence into a cohesive framework. Our algorithm skillfully navigates the complexities inherent in these networks, enhancing traditional analysis methods by incorporating algebraic connectivity. This approach allows for the simultaneous consideration of multiple layers and their intricate interactions. Through extensive validation against established centrality measures and application to a variety of real-world datasets, LayerPlexRank has proven its ability to deliver nuanced insights into the structure and dynamics of complex networks. Looking ahead, there is potential to further refine the algorithm by expanding the centrality calculations to encompass interconnected multilayered networks with explicit inter-layer connections. This development could significantly push the boundaries of current research in network science, opening new avenues for exploring the depth and breadth of network interactions.

### v2

@@ Abstract.
Centrality; Influence; Algebraic Connectivity; Multiplex Networks
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USAbooktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAdoi: 10.1145/3627673.3679950isbn: 979-8-4007-0436-9/24/10ccs: Mathematics of computing Graph algorithmsccs: Mathematics of computing Network flowsccs: Information systems Content rankingccs: Information systems Learning to rankccs: Information systems Similarity measuresccs: Theory of computation Network flows
@@ 3.3. Robustness by Designed Cross-validation for Multiplex Networks
In experiments, we demonstrate robustness using a specifically designed Leave-one-out Cross Validation (LOOCV) method (Sammut and Webb, 2010). LOOCV can be considered a specific application of cross-validation, which splits the original $n$-node dataset into an $n-1$ node training dataset and a $1$-node validation dataset in each iteration. The total number of validation times is $n$, indicating that all nodes have been left out once and used as a validation dataset. However, LOOCV focuses on multiple evaluations of the remaining training set using the split validation set, while our algorithm does not predict unknown data from existing data. Therefore, we use Jackknife resampling (Berger and Rao, 2006), which omits a node and its edges in each iteration, complementing LOOCV to robustly aggregate results. This method highlights our algorithm’s resilience to disruptions.
@@ 4. Conclusions
LayerPlexRank advances multiplex network analysis by integrating node centrality and layer influence into a cohesive framework. Our algorithm effectively navigates the complexities of these networks, enhancing traditional methods with algebraic connectivity. This allows for the simultaneous consideration of multiple layers and their interactions. Extensive validation against established centrality measures and application to various real-world datasets demonstrate LayerPlexRank’s ability to deliver nuanced insights into complex network structures and dynamics. Future refinement could expand centrality calculations to interconnected multilayered networks with explicit inter-layer connections, potentially pushing the boundaries of network science research.
Acknowledgements.
This work is supported by Australian Research Council Discovery Early Career Award (project number DE240101049).